OpenStudy (anonymous):

if i have 8 over 10 and 6 over 10 adding it and simplify what is the answer? please help

OpenStudy (anonymous):

8/10 = 4/5 6/10 = 3/5 4/5+3/5 = 8/5

OpenStudy (gw2011):

The answer is actually 7/5 = 1 2/5 = 1.4